# Zero shot VLMs for hate meme detection: Are we there yet?

### v2

@@ Abstract
Multimedia content on social media is rapidly evolving, with memes gaining prominence as a distinctive form. Unfortunately, some malicious users exploit memes to target individuals or vulnerable communities, making it imperative to identify and address such instances of hateful memes. Extensive research has been conducted to address this issue by developing hate meme detection models. However, a notable limitation of traditional machine/deep learning models is the requirement for labeled datasets for accurate classification. Recently, the research community has witnessed the emergence of several visual language models that have exhibited outstanding performance across various tasks. In this study, we aim to investigate the efficacy of these visual language models in handling intricate tasks such as hate meme detection. We use various prompt settings to focus on zero-shot classification of hateful/harmful memes. Through our analysis, we observe that large VLMs are still vulnerable in zero-shot hate meme detection. Warning: Contains potentially offensive memes.
1 Introduction
@@
Several large vision language models (VLMs) have recently become available to the public. These models exhibit impressive performance across various tasks, including sentiment analysis Kheiri and Karimi (2023), visual question answering Lan et al. (2023), and so on. However, the existing literature falls short in addressing how VLMs perform precisely in the context of hate meme detection Van and Wu (2023) and particularly in a zero-shot setting. The urgency for such systems stem from the exponential growth in multi-modal content on social media platforms with malevolent individuals severely exploiting memes as a tool to target various communities and propagate hate Gomez et al. (2020). While manual moderation is nearly impossible, traditional machine learning models can also be not extensively trained for automatic moderation due to the severe lack of labeled hateful memes datasets that are diverse in terms of language, target groups and social setting. This gap in research underscores the need to explore and evaluate the effectiveness of zero-shot VLMs for identifying and mitigating the spread of such content in memes. Note that the zero-shot setting is important here since curating labeled hateful meme datasets that are socially, culturally and target-wise diverse is extremely difficult.
In this paper, for the first time, we systematically employ various prompt strategies and input instructions to assess the ‘power’ of well-known open source VLMs, including IDEFICS Laurençon et al. (2023), LLaVA-1.5 Liu et al. (2023), and InstructBLIP Dai et al. (2023) in detecting hateful memes in a fully zero-shot setting. Note that we purposefully choose open source models since they do not come with a huge monetary cost unlike their commercial counterparts. We evaluate the outputs of these models for four well-known datasets covering hateful, misogynistic, and harmful memes. The central contributions of this paper are as follows.
@@
(i) Systematic evaluation of classification capability of VLMs: We systematically study the effect of prompt strategies that we use to query these models to understand their strengths and vulnerabilities. In total we investigate as many as 32 prompts (8 prompt variations across 4 datasets) for each model. This is unlike what is typically done in a majority of studies where the model is queried using one or two prompt variants at most thus limiting the true potential of prompt engineering. Our prompts can be broadly categorized into the following types based on the input and output patterns: input variants can comprise (a) vanilla input, (b) input along with the definition of what is hateful/misogynistic/harmful, (c) input along with OCR text, (d) input along with definition and OCR text; output variants can be (a) vanilla output, (b) output along with an explanation. We observe that prompt variants that are most successful in eliciting correct responses vary across models and datasets; nevertheless, in many of the cases OCR text alone or OCR text with definition works well.
@@
Overall, our study has a far larger scope than the standard objective of identifying the best all-purpose VLMs. It strives to rather choose the best prompt variant across different models using a thorough and principled prompt engineering approach. Further it lays a foundation to identify interpretable typological categories of hateful memes that the VLMs are most vulnerable to. These induced topics can be used to improve the performance of VLMs by implementing safety guardrails without fine-tuning the models repeatedly which typically comes with a huge compute cost.
2 Related works
Hate meme detection: A growing body of research in recent years focused on hate meme detection Gomez et al. (2020); Kirk et al. (2021); Shang et al. (2021). Several datasets and models have been developed, encompassing various dimensions, including hateful content detection Kiela et al. (2020), misogyny detection Fersini et al. (2022), cyberbullying detection Maity et al. (2022), harmful meme detection Pramanick et al. (2021a, b), and many more Chandra et al. (2021); Lin et al. (2024) including other languages Das and Mukherjee (2023).
Vision language models: IDEFICS Laurençon et al. (2023), LLaVA-1.5 Liu et al. (2023), InstructBLIP Dai et al. (2023), Flamingo Alayrac et al. (2022), and OpenAI GPT-4 AI are popular vision language models widely used for tasks like sentiment analysis Kheiri and Karimi (2023), question answering Lan et al. (2023), and chatbot creation AlZu’bi et al. (2024), hate meme detection Van and Wu (2023). However, exploring hate meme detection using VLMs is limited Van and Wu (2023); Lin et al. (2024), particularly in the context of different prompt scenarios, different model setups and thorough interpretation of results.
3 Datasets and metrics
Datasets: This section introduces the four datasets we have utilized to explore the performance of Vision-Language Models (VLMs). These datasets cover three types of memes: hate, misogyny, or harmful content (see Table 1 for details).
@@ 3 Datasets and metrics
Table 1: Label distribution for each dataset.
@@
(1) Facebook hateful memes (FHM): The FHM dataset introduced by Facebook AI Kiela et al. (2020) is a collection of memes designed to help researchers develop tools for identifying and removing hateful content online. The dataset contains more than 10K memes labeled hateful and not-hateful, covering various targets, including race, ethnicity, religion, gender, sexual orientation, and disability. We use a random sample of 500 memes [^1] Note that for this dataset the test set was removed by the authors after the competition. We have therefore used the validation set to sample our data points. in order to test the VLMs in the zero-shot setting.
@@
(2) Multimedia automatic misogyny identification (MAMI): The MAMI Fersini et al. (2022) dataset was created and shared as part of SemEval-2022 Task 5. Unlike the FHM dataset, the MAMI dataset focuses on identifying misogyny in online memes. The dataset contains 11K memes, of which 1K memes are in the test set, and we conduct all our experiments considering only the test set. Each meme has a binary label – ‘misogynous’ or ‘not misogynous’ – which we use for our experiments.
@@
(3) Harmful memes: ‘Harmful’ is a more general term compared to ‘offensive’ and ‘hateful’. While an offensive or hateful meme is harmful, not all harmful memes are necessarily hateful or offensive. We utilize the HARM-P Pramanick et al. (2021b) (related to US politics) and HARM-C Pramanick et al. (2021a) (related to COVID-19) datasets for our experiments. Both datasets contain more than 3.5K memes. For our study, we only consider the test sets. The original labels of both datasets have three classification labels: not harmful, somewhat harmful, and very harmful. To maintain consistency with our binary classification experiments, we have merged somewhat harmful and very harmful into a single category labeled as harmful.
@@
Metrics: As we perform binary classification tasks, we measure the models’ performance using accuracy, macro F1 score, and area under the ROC curve metrics.
4 Models
We ran our experiments on a total of five different models. All models are open source large VLMs. Due to resource constraints as well as to make a fair comparison, we ran our experiments using 8-bit quantization Liu et al. (2021) for all the VLMs.
@@ 4 Models
IDEFICS: IDEFICS Laurençon et al. (2023) which closely follows the architecture of Flamingo, is trained on open source datasets like OBELICS and LAION. It combines two frozen uni-modal backbones which are, LLaMA as the language model and OpenClip as the vision encoder. We used instruction fine-tuned IDEFICS 9B model with the checkpoint HuggingFaceM4/idefics-9b-instruct for our experiments.
@@
LLaVA-1.5: LLaVA-1.5 Liu et al. (2023) is an enhanced version of LLaVA. LLaVA combines LLaMA/Vicuna as the language model and CLIP as the vision encoder. Compared to LLaVA, LLaVA-1.5 has enhanced capabilities due to the addition of an MLP vision-language connector and integration of academic task-oriented data. We have used two different LLaVA-1.5 models with 7B and 13B parameters. The checkpoints of these models are llava-hf/llava-1.5-7b-hf and llava-hf/llava-1.5-13b-hf.
@@
InstructBLIP: InstructBLIP Dai et al. (2023) is an instruction fine-tuned model that uses the same architecture as BLIP-2 with a small but significant difference. It uses frozen Flan-T5/Vicuna as the language model and a vision transformer as the image encoder. Extending BLIP-2, InstructBLIP proposes an instruction-aware Q-Former module. As additional inputs, the model takes instruction text tokens which interacts with the query embeddings via the self-attention layer of the Q-Former. We have used two different InstructBLIP models with Vicuna 7B and Flan-T5-xl as backbone language models. The checkpoints of these models are Salesforce/instructblip-vicuna-7b and Salesforce/instructblip-flan-t5-xl respectively.
5 Prompts
@@
Vanilla input: Following Roy et al. (2023), we use a prompt template to instruct the model to classify the given meme into a label from a predefined list_of_labels. However, in our scenario, the list_of_labels is only restricted to binary labels. In addition, we supply two example_outputs (one label per line for positive and negative samples) to assist the models in generating appropriate answers. In our case, ‘positive’ denotes content deemed hateful, misogynistic, or harmful based on the dataset passed to the model.
(+) Definition input: For vanilla prompts, we assumed that VLMs are to some extent aware of the labels for classifying the input image. Here, we take a step further and add the definition of the labels as an additional context to the VLMs. Our intuition was similar to Roy et al. (2023), i.e., the definition can help the VLMs understand the classification tasks better. We picked and added one line of definition from the corresponding dataset for all list_of_labels (positive and negative in our case). We provide definitions of the labels for each dataset in Appendix A.
(+) OCR input: In a meme, multi-modality, i.e., embedded text and image play very crucial role in the classification task, similar to the works Pramanick et al. (2021a); Das and Mukherjee (2023). We therefore add ocr_extracted_text in the vanilla prompt. Our intuition was that the models would further be better in understanding the contexts with this addition and would be more successful in classifying the input image meme as per the list_of_labels. We provide the ocr_extracted_text enclosed within three back-ticks for the model to easily distinguish it from other texts in the prompt.
@@
(+) Explanation output: Adding to the above case of vanilla output, we prompt the model to further explain the raionale (within 30 words) based on which it made a prediction.
@@
6 Experimental setup
For all the models, we use a batch size of 1. We manually tune the temperature values and set them to 1.0 for the IDEFICS, LLaVA-1.5 7B and 13B models, and 0.8 for the InstructBLIP models. The temperature parameter controls how random the generated output would be. However, with lower temperatures, we observed inferior performance of these models. As noted earlier, we experiment with eight different prompts on four datasets, studying them across five models. In short, we run 32 prompts per model and 160 prompts across all five models. All the models are coded in Python using the PyTorch library. We utilize 2xT4 GPUs from Kaggle, providing a total of 15GB memory on each GPU with a usage limit of 30hrs/week. Further setup details are provided in Appendix C. We present the detailed results in the following section.
7 Results
[…]
Strategies FHM MAMI HARM-C HARM-P
in out acc mf1 auc acc mf1 auc acc mf1 auc acc mf1 auc

### v3

# Exploring the Limits of Zero Shot Vision Language Models for Hate Meme Detection: The Vulnerabilities and their Interpretations

@@ Abstract
There is a rapid increase in the use of multimedia content in current social media platforms. One of the highly popular forms of such multimedia content are memes. While memes have been primarily invented to promote funny and buoyant discussions, malevolent users exploit memes to target individuals or vulnerable communities, making it imperative to identify and address such instances of hateful memes. Thus social media platforms are in dire need for active moderation of such harmful content. While manual moderation is extremely difficult due to the scale of such content, automatic moderation is challenged by the need of good quality annotated data to train hate meme detection algorithms. This makes a perfect pretext for exploring the power of modern day vision language models (VLMs) that have exhibited outstanding performance across various tasks. In this paper we study the effectiveness of VLMs in handling intricate tasks such as hate meme detection in a completely zero-shot setting so that there is no dependency on annotated data for the task. We perform thorough prompt engineering and query state-of-the-art VLMs using various prompt types to detect hateful/harmful memes. We further interpret the misclassification cases using a novel superpixel based occlusion method. Finally we show that these misclassifications can be neatly arranged into a typology of error classes the knowledge of which should enable the design of better safety guardrails in future [^1] Accepted at AAAI (ICWSM) 2025. Code and other relevant sources are available online [^2] https://github.com/hate-alert/HateVLMs.
Warning: Contains potentially offensive content.
Introduction
@@
Several large vision language models (VLMs) have recently become available to the public (Kheiri and Karimi 2023; Lan et al. 2023). A pertinent question is how VLMs perform precisely in the context of hate meme detection (Plaza-del arco, Nozza, and Hovy 2023; Van and Wu 2023) and particularly in a zero-shot setting. The urgency for such systems stem from the exponential growth in multi-modal content on social media platforms and companies like Meta releasing statements of using LLMs/VLMs for content moderation [^3] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/. This choice is justified since while manual moderation is nearly impossible, traditional machine learning models can also be not extensively trained for automatic moderation due to the severe lack of labeled hateful memes datasets that are diverse in terms of language, target groups and social setting. This gap in research underscores the need to explore and evaluate the effectiveness of zero-shot VLMs for identifying and mitigating the spread of such content in memes. Note that the zero-shot setting is important here since curating labeled hateful meme datasets that are socially, culturally and target-wise diverse is extremely difficult.
In this paper, for the first time, we systematically employ various prompt strategies and input instructions to assess the ‘power’ of well-known VLMs, including IDEFICS (Laurençon et al. 2023), LLaVA-1.5 (Liu et al. 2023), InstructBLIP (Dai et al. 2023) and GPT-4o (OpenAI 2024) in detecting hateful memes in a fully zero-shot setting. Our work is motivated by the recent literature on zero-shot evaluation of LLMs (Plaza-del arco, Nozza, and Hovy 2023; Roy et al. 2023; Saha et al. 2024). Not only there is an extreme shortage of quality datasets for hateful meme detection but also as we shall show, many memes (due to their implicit nature and nuanced presentation) routinely get incorrectly annotated by humans across benchmark datasets. Thus to make these models usable for moderation in online platforms, it is important to understand their zero-shot capabilities carefully gauging the risks they pose on unseen data so that necessary fixes can be implemented in future. Consequently, our main goal here is to perform interpretability analysis to understand the vulnerabilities faced by these models in a completely zero-shot setup. While fine-tuning the model might improve the results it might also limit the generalizability of the model to the type of data it is being finetuned on. Further, fine-tuning with incorrectly annotated data might result in more harm than good.
We evaluate the outputs of these models for four well-known datasets covering hateful, misogynistic, and harmful memes in English and two datasets covering hateful, and offensive memes in Bengali and HinGlish [^4] Hinglish is a blend of Hindi and English, where speakers mix words, phrases, or sentences from both languages in conversation. language respectively. The central contributions of this paper are as follows.
@@
(i) Systematic evaluation of classification capability of VLMs: We systematically study the effect of prompt strategies that we use to query these models to understand their strengths and vulnerabilities. In total we investigate as many as 48 prompts (8 prompt variations across 6 datasets) for each model. This is unlike what is typically done in a majority of studies where the model is queried using one or two prompt variants at most thus limiting the true potential of prompt engineering. Our prompts can be broadly categorized into the following types based on the input and output patterns: input variants can comprise (a) vanilla input, (b) input along with the definition of what is hateful/misogynistic/harmful/offensive, (c) input along with OCR [^5] OCR - Optical Character Recognition: In memes, this tool is used to extract the embedded textual content. In our work, we use the OCR extracted text already present in the datasets. text, (d) input along with definition and OCR text; output variants can be (a) vanilla output, (b) output along with an explanation. We observe that prompt variants that are most successful in eliciting correct responses vary across models and datasets; nevertheless, in many of the cases OCR text alone or OCR text with definition works well.
@@
Overall, our study has a far larger scope than the standard objective of identifying the best all-purpose VLMs. It strives to rather choose the best prompt variant across different models using a thorough and principled prompt engineering approach. Further it lays a foundation to identify interpretable typological categories of hateful memes that the VLMs are most vulnerable to. These induced topics can be used for actionable evaluation (Vilar et al. 2006) to improve the performance of VLMs by implementing safety guardrails without fine-tuning the models repeatedly which typically comes with a huge compute cost. We outline the summary and flow of the paper in Figure 1.
Related works
Figure 1: Pipeline: A concise summary and flow of the evaluation and analysis carried out in this work.
Hate meme detection: A growing body of research in recent years focused on hate meme detection (Gomez et al. 2020; Kirk et al. 2021; Shang et al. 2021). Several datasets and models have been developed, encompassing various dimensions, including hateful content detection (Kiela et al. 2020), misogyny detection (Fersini et al. 2022), cyberbullying detection (Maity et al. 2022), harmful meme detection (Pramanick et al. 2021a, b), and many more (Chandra et al. 2021; Lin et al. 2024) including other languages (Das and Mukherjee 2023; Hossain et al. 2024; Kumari et al. 2024).
Vision language models: IDEFICS (Laurençon et al. 2023), LLaVA-1.5 (Liu et al. 2023), InstructBLIP (Dai et al. 2023), Flamingo (Alayrac et al. 2022), and GPT-4o (OpenAI 2024) are popular vision language models widely used for tasks like sentiment analysis (Kheiri and Karimi 2023), question answering (Lan et al. 2023), and hate meme detection (Van and Wu 2023). However, exploring hate meme detection using VLMs is limited (Van and Wu 2023; Lin et al. 2024), particularly in the context of different prompt scenarios, different model setups & languages and thorough interpretation of results.
Interpretability: With the advent of deep learning, several efforts are made to explain the working of the model’s predictions (Selvaraju et al. 2017; Gyevnar, Ferguson, and Schafer 2023; Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017). Recently, there have also been efforts to evaluate vision-language multimodal systems (Chen et al. 2023; Zhao et al. 2022; Lei et al. 2023). However, memes are a special type of image where the actual context is inherently decided by the combination of image and the embedded text within the meme. There are a handful of works present on understanding memes as well (Deshpande and Mani 2021; Hee, Lee, and Chong 2022; Hee, Chong, and Lee 2023). However none of these works present the study of important regions in the meme that lead to model’s prediction with respect to an input prompt given to the VLMs. More importantly, to the best of our knowledge, there are no works presently that can interpret the predictions of closed source VLMs like GPT-4o. In this work, we propose a novel model agnostic superpixel based approach to effectively understand the important regions in the meme leading to model’s prediction corresponding to an input prompt.
Datasets and metrics
Datasets: This section introduces the six datasets we have utilized to explore the performance of Vision-Language Models (VLMs). These datasets cover four types of memes: hate, misogyny, harmful, or offensive content spanning three languages: English, Bengali & HinGlish. (see Table 1 for details).
@@
BHM Hateful 266 711
Not hateful 445
HinGlish Offensive 250 500
Not offensive 250
Table 1: Datasets: Label distribution for each dataset.
@@
(1) Facebook hateful memes (FHM): The FHM dataset introduced by Facebook AI (Kiela et al. 2020) is a collection of memes designed to help researchers develop tools for identifying and removing hateful content online. The dataset contains more than 10K memes labeled hateful and not-hateful, covering various targets, including race, ethnicity, religion, gender, sexual orientation, and disability. We use a random sample of 500 memes [^6] Note that for this dataset the test set was removed by the authors after the competition; therefore we used the dev split. in order to test the VLMs in the zero-shot setting.
@@
(2) Multimedia automatic misogyny identification (MAMI): The MAMI (Fersini et al. 2022) dataset was created and shared as part of SemEval-2022 Task 5. Unlike the FHM dataset, the MAMI dataset focuses on identifying misogyny in online memes. The dataset contains 11K memes, of which 1K memes are in the test set, and we conduct all our experiments considering only the test set. Each meme has a binary label – misogynous or not misogynous – which we use for our experiments.
@@
(3) Harmful memes: ‘Harmful’ is a more general term compared to ‘offensive’ and ‘hateful’. While an offensive or hateful meme is harmful, not all harmful memes are necessarily hateful or offensive. We utilize the HARM-P (Pramanick et al. 2021b) (related to US politics) and HARM-C (Pramanick et al. 2021a) (related to COVID-19) datasets for our experiments. Both datasets contain more than 3.5K memes. For our study, we only consider the test sets. The original labels of both datasets have three classification labels: not harmful, somewhat harmful, and very harmful. To maintain consistency with our binary classification experiments, we have merged somewhat harmful and very harmful into a single category labeled as harmful.
(4) Bangla hateful memes: The BHM dataset was proposed in a recent work by (Hossain et al. 2024) and contains over 7K memes in Bengali language. These memes are labelled as hateful or not-hateful and are specifically prevalent among Bengali speakers in India & Bangladesh, covering a variety of discussion topics. We use the test split with 711 memes for our experiments.
(5) HinGlish offensive memes: HinGlish is a language commonly spoken within Indian subcontinent and is a mix of Hindi with English. A recent work by (Kumari et al. 2024) introduced a dataset in HinGlish which contains offensive or not-offensive memes. Although the paper mentions around 7k samples, just over 2k memes have been made public [^7] https://github.com/Gitanjali1801/CM˙MEMES. We reached out to the authors of the paper for the complete dataset but without any success. The memes surround the activities which are prevalent within Indian subcontinent. Since the full dataset has not been made public, we randomly sample 250 memes from each of the offensive and not offensive classes.
@@
Metrics: As we perform binary classification tasks, we measure the models’ performance using accuracy, and macro-F1 score as metrics.
Models
We ran our experiments on a total of six different models.
@@
IDEFICS: IDEFICS (Laurençon et al. 2023) which closely follows the architecture of Flamingo, is trained on open source datasets like OBELICS and LAION. It combines two frozen uni-modal backbones which are, LLaMA as the language model and OpenClip as the vision encoder. We used instruction fine-tuned IDEFICS 9B model with the checkpoint HuggingFaceM4/idefics-9b-instruct for our experiments.
@@
LLaVA-1.5: LLaVA-1.5 (Liu et al. 2023) is an enhanced version of LLaVA. LLaVA combines LLaMA/Vicuna as the language model and CLIP as the vision encoder. Compared to LLaVA, LLaVA-1.5 has enhanced capabilities due to the addition of an MLP vision-language connector and integration of academic task-oriented data. We have used two different LLaVA-1.5 models with 7B and 13B parameters. The checkpoints of these models are llava-hf/llava-1.5-7b-hf and llava-hf/llava-1.5-13b-hf.
@@
InstructBLIP: InstructBLIP (Dai et al. 2023) is an instruction fine-tuned model that uses the same architecture as BLIP-2 with a small but significant difference. It uses frozen Flan-T5/Vicuna as the language model and a vision transformer as the image encoder. Extending BLIP-2, InstructBLIP proposes an instruction-aware Q-Former module. As additional inputs, the model takes instruction text tokens which interacts with the query embeddings via the self-attention layer of the Q-Former. We have used two different InstructBLIP models with Vicuna 7B and Flan-T5-xl as backbone language models. The checkpoints of these models are Salesforce/instructblip-vicuna-7b and Salesforce/instructblip-flan-t5-xl respectively.
GPT-4o: GPT-4o (OpenAI 2024) (‘o’- omni) is a model released by OpenAI that takes input an arbitrary combination of image or text and generates textual output. Compared to previous OpenAI models, GPT-4o is comparatively better at vision understanding. In our work, this is the only closed source model which we have considered apart from the five open source models. We would like to thank Microsoft Azure for providing us with the credits to use their OpenAI services. We used the provided APIs to infer GPT-4o from their documentation to run our experiments [^8] https://learn.microsoft.com/en-us/azure/ai-services/openai/ and for our case, we used a combination of single image with text as input to remain consistent with the zero-shot setup.
Baseline: Our work focuses on the evaluation of zero-shot capability of VLMs. To have a fair comparison with a very recent prior work presented in (Cao et al. 2023), we consider two different baselines: (i) complete zero-shot assessment, and (ii) fine-tuned only on FHM (a very diverse dataset) for two epochs (instead of ten; to be very near to the zero-shot setup) and evaluation across all six datasets. These setups not only allow for fair evaluation, but also help in assessing the generalization capability, which is often ignored. Two samples (one negative and one positive) are also provided with the test sample in the prompt, as noted in the mentioned paper; hence we present results as mean and standard deviation across three different seeds for each setup.
Prompts
@@
Vanilla input: Following (Roy et al. 2023), we use a prompt template to instruct the model to classify the given meme into a label from a predefined list_of_labels. However, in our scenario, the list_of_labels is only restricted to binary labels. In addition, we supply two example_outputs (one label per line for positive and negative samples) to assist the models in generating appropriate answers. In our case, ‘positive’ denotes content deemed hateful, misogynistic, harmful, or offensive based on the dataset passed to the model.
(+) Definition input: For vanilla prompts, we assumed that VLMs are to some extent aware of the labels for classifying the input image. Here, we take a step further and add the definition of labels as an additional context to the VLMs. Our intuition was similar to (Roy et al. 2023), i.e., the definition can help the VLMs understand the classification tasks better. We picked and added one line of definition from the corresponding dataset for all list_of_labels (positive and negative in our case). We provide definitions of the labels for each dataset in AppendixA.
(+) OCR input: In a meme, multi-modality, i.e., embedded text and image play very crucial role in classification, similar to the works (Pramanick et al. 2021a; Das and Mukherjee 2023). We therefore add ocr_extracted_text in the vanilla prompt. Our intuition was that the models would further be better in understanding the contexts with this addition and would be more successful in classifying the input meme as per the list_of_labels. We provide the ocr_extracted_text enclosed within three back-ticks for the model to distinguish it from other texts in the prompt.
@@
(+) Explanation output: Adding to the above case of vanilla output, we prompt the model to further explain the rationale (within 30 words) based on which it made a prediction.
@@
The definition of hate speech, the explanation as to why a post is hateful, and the OCR text (often indicating the victim community) are a hate meme’s three most important components. We believe that prompts should at least include information about these components, as has also been noted in (Roy et al. 2023). We agree that there can be other variants but at meta level the above three units should always be considered while framing a prompt for hate meme detection.
Experimental setup
For all the models, we use a batch size of 1. We manually tune the temperature values and set them to 1.0 for the IDEFICS, LLaVA-1.5 7B & 13B, GPT-4o models, and 0.8 for the InstructBLIP models. The temperature parameter controls how random the generated output would be. However, with lower temperatures, we observed inferior performance of these models. As noted earlier, we experiment with eight different prompts on six datasets, studying them across six models. In short, we run 48 prompts per model and 288 prompts across all six models. All the open source models are coded in Python using the PyTorch library while we have no information about GPT-4o since it is a proprietary model. For open source models, we utilize 2xT4 GPUs from Kaggle, providing a total of 15GB memory on each GPU with a usage limit of 30hrs/week We provide further setup details for these models in AppendixC. We present the detailed results in the following section.
Results
Strategies FHM MAMI HARM-C HARM-P BHM HinGlish
in out acc mf1 acc mf1 acc mf1 acc mf1 acc mf1 acc mf1